Simplify using factoring and cancellation
(y^3-19y-30)/(y+2)

Answer by fractalier(2101)

Though it is not easy to see,
(y^3 - 19y - 30) / (y + 2) =
(y + 2)(y^2 - 2y - 15)/ (y + 2) =
y^2 - 2y - 15